OpenStudy (anonymous):

Find the limit of (1-3/x)^2x as x approaches positive infinity

OpenStudy (anonymous):

Is anyone here taking up calculus?

sam (.sam.):

Combine 1-(3)/(x) into a single expression \[\lim_{x \rightarrow \infty}(\frac{x-3}{x})\] The value of \[\lim_{x \rightarrow \infty}(\frac{x-3}{x})\] is 1

OpenStudy (zarkon):

have you seen \[\lim_{n\to\infty}\left(1+\frac{x}{n}\right)^{n}=e^x\]

OpenStudy (anonymous):

Why is the answer 1?

OpenStudy (anonymous):

Mr. Zarkon is good so you can do too write like this [(1 + -3/x)^x]^2

OpenStudy (anonymous):

by the way that equation is still raised to 2x

sam (.sam.):

ignore my answer follow zarkon's one @Naomi22 yes i didn't notice there's power 2x

OpenStudy (zarkon):

I prefer Lord Zarkon, King Zarkon or Dr. Zarkon... as appose to Mr. Zarkon :)

OpenStudy (anonymous):

Ok yes sir.. thank you.

OpenStudy (anonymous):

alright. so how can I write this in terms of e

OpenStudy (zarkon):

combine what WonHungLo and I wrote

OpenStudy (anonymous):

e^[(1-3/x)^2x] like that?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

oh yeah where did the limit go?

sam (.sam.):

you e^(limit)

OpenStudy (anonymous):

oh ok

OpenStudy (zarkon):

\[\lim_{x\to\infty}\left(1+\frac{a}{x}\right)^{b x}=e^{ab}\]

OpenStudy (anonymous):

is this right? e^lim[(1-3/x)^2x]

OpenStudy (zarkon):

no

sam (.sam.):

nvm ,use zarkon's way

OpenStudy (anonymous):

I'll just directly copy zarkon's equation on paper then plug in my values. It's easier that way

OpenStudy (zarkon):

what do you get for your final answer?

OpenStudy (anonymous):

could you solve it by the way?

OpenStudy (anonymous):

wait I'm getting 0 x infinity

OpenStudy (zarkon):

I can...can you?...what do you get? use the formula I gave you

OpenStudy (anonymous):

we're told to use L'hopital's rule on this one

OpenStudy (anonymous):

but using your formula it's e^-6

OpenStudy (zarkon):

yes...that is the answer

OpenStudy (anonymous):

great thanks. but I still have to figure out a way to do this applying L'hopital's rule

OpenStudy (zarkon):

if you want to calculate it with out the formula then \[\lim_{x\to\infty}\left(1-\frac{3}{x}\right)^{2x}\] let \[y=\left(1-\frac{3}{x}\right)^{2x}\] then \[\ln(y)=2x\ln\left(1-\frac{3}{x}\right)\] which can be written as \[\ln(y)=2\frac{\ln\left(1-\frac{3}{x}\right)}{\frac{1}{x}}\] now take the limit using L'Hospitals rule

OpenStudy (anonymous):

alright thank you very much Lord Zarkon. Your name fits you just fine :D

OpenStudy (zarkon):

no problem